# Does provable absence of barren plateaus imply classical simulability? Or, why we need to rethink variational quantum computing

### v2

@@ III What leads to absence of barren plateaus?
If, however, the unitary $U(\bm{\theta})$ possesses additional structure, such as respecting some symmetry, then the loss can inherit this structure and potentially avoid the aforementioned issues. In particular, let us consider the adjoint action of $U(\bm{\theta})$ over the operator space $\mathcal{B}$, and let us analyze if it leads to either subspaces, or, effective subspaces. That is, given some $\bm{\theta}$ and operator $P$ in an appropriate orthogonal basis of $\mathcal{B}$ (such as a Pauli operator), we want to know: Where can $U(\bm{\theta})PU^{\dagger}(\bm{\theta})$ go? [^2] Note that one can, and should, also ask the question: Where can $U^{\dagger}(\bm{\theta})PU(\bm{\theta})$ go? But for simplicity of notation we will consider the one in the main text.
Mathematically, this question can be answered by computing the inner products $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ for the rest of the basis elements $P_{j}$. For instance, if the inner product is non-zero for all $P_{j}$, then we know that the unitary can transform $P$ into an operator that can spread out and reach operators across all of $\mathcal{B}$. On the other hand, it could happen that the adjoint action of the unitary can only reach certain operators in $\mathcal{B}$. For instance, as shown in Fig. 3(a), the adjoint action of $U(\bm{\theta})$ can lead to well-defined subspaces in $\mathcal{B}$ such that $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ is only non-zero for operators in some $\mathcal{B}_{P}\subset\mathcal{B}$. This case can arise for instance in circuits with small Lie algebraic modules [20, 64, 86, 87, 50, 51, 52] or in shallow-depth hardware efficient ansätze [22, 23, 24, 25].
A second case of interest arises when we have that $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ is non-zero for a wide range of operators, but is large only for operators in a given subspace $\mathcal{B}_{P}$. This case is depicted schematically in Fig. 3(b) and we will say that given $P$, the adjoint action of $U(\bm{\theta})$ leads to ‘effective’ subspaces, rather than proper ones. We note that effective subspaces can either arise for all $\bm{\theta}$ or with high probability for $\bm{\theta}\sim\mathcal{P}$. In the latter case, for most $\theta$ values for the $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ are large only within a small subspace, but for some low probability $\bm{\theta}$ values, large overlaps outside this subspace could be observed. Effective subspaces appear, with high probability for $\bm{\theta}\sim\mathcal{P}$, for quantum convolutional neural networks [88, 13], or in circuits with small angle initialization [37, 53, 58].
@@
Shallow hardware efficient ansatz [22, 23, 24, 25, 27, 28, 29] Acting on $\mathcal{O}(\log(n))$ neighboring qubits (P) Local $O$, area law $\rho$
Quantum convolutional neural network [13] Acting on $\mathcal{O}(1)$ qubits (E) Local $O$, area law $\rho$
@@
[…]
Small angle initialization [37, 53, 58] Acting on $\mathcal{O}(1)$ qubits (E) Local $O$, area law $\rho$
[…]
@@
Table 1: Problem classes in $\overline{{\rm BP}}$ and also in $\mathcal{C}_{{\rm polySub}}$. In this table, we present the problems considered in our analysis. In all cases, we also provide the references where absence of barren plateaus was proven, as well as the relevant polynomially large subspace $\mathcal{B}_{\lambda}$ (indicating whether it is proper (P) or effective (E)), and the states and measurement operators for which $\mathcal{C}$ is in $\overline{{\rm BP}}$. We refer the reader to Table 2 and the Appendix for additional details on how the classical simulation is performed. Here ‘proj.’ is used as a short-hand for projector, while MPS stands for matrix product states [89].
To showcase the importance of having measurement operators and initial states which are ‘well-aligned’ with the polynomial subspace, let us consider again the class of problems with shallow hardware efficient ansätze. We already saw that if $O$ is local (or contains relevant local terms), then $\mathcal{C}_{{\rm shallowHEA}}\subset\mathcal{C}_{{\rm polySub}}$. Thus, to know whether the loss will concentrate, we need to study the Hilbert-Schmidt norm of the projected $\rho$ into the subsets of $2L\in\mathcal{O}(\log(n))$ adjacent qubits in the backwards light cone of the local measurement. For instance, it is not hard to check that if $\rho$ is pure and follows a volume law of entanglement [40], then $|\left\langle\rho_{\lambda},\rho_{\lambda}\right\rangle|$ will be exponentially close to zero, and the loss will be exponentially concentrated. On the other hand, if $\rho$ is pure and satisfies an area law of entanglement [90, 40], then the previous inner product will be, at most, polynomially vanishing. Putting these realizations together we see that
@@
Standard provably barren plateau-free architectures live in classically identifiable polynomial subspaces.
For all standard problem classes $\mathcal{C}$ such that $\mathcal{C}\subset\overline{{\rm BP}}$, we have that
@@
By ‘standard problems’ we here refer to the conventional variational quantum architectures which have been proven to be in $\overline{{\rm BP}}$. In Table 1 we present a non-exhaustive list of such architectures. Therein, we report the relevant polynomial subspace, indicating whether it is proper or effective, as well as the conditions on the measurement operator and initial state necessary for absence of barren plateaus. We emphasize that in many cases, it was the proof of absence of barren plateau itself which allowed us to determine the reported information. The majority of strategies studied, including shallow hardware efficient ansatz and highly symmetrized models, lead to proper subspaces for all $\bm{\theta}$. The most important examples of models where one lives in effective subspaces (with high probability for $\bm{\theta}\sim\mathcal{P}$) are quantum convolutional neural networks and small angle initialization strategies.
@@ IV Connection between absence of barren plateaus and simulability
$\mathcal{C}_{{\rm shallowHEA}}\subset{\rm CSIM}\,,\quad\text{or}\quad\mathcal{C}_{{\rm shallowHEA}}\subset{\rm CSIM}_{{\rm QE}}\,,$ (12)
and it is clear that the loss can be estimated without ever needing to use a quantum computer to run the parametrized quantum circuit. Hence, no problem based on shallow hardware efficient ansätze with a local measurement can be outside of ${\rm CSIM}_{{\rm QE}}$. We note that this result was originally reported in Ref. [28]. [^4] It is worth additionally highlighting that, given the simplicity of the requisite measurement procedures here, one could lift the restriction that $\rho$ is preparable by a poly depth quantum circuit and also consider quantum states that result from quantum experiments running in polynomial time, e.g., from analog or non-unitary processes. This makes the quantum-enhanced classical simulation method more flexible than directly implementing the parameterized quantum circuit on hardware [91].
@@
Shallow hardware efficient ansatz Pauli classical shadows [78] Reduced $U(\bm{\theta})$ For all $\bm{\theta}$
Quantum convolutional neural network [^5] In the Appendix we also show that certain parameter restricted quantum convolutional neural networks can be simulated for all $\bm{\theta}$. Pauli classical shadows Tensor Networks For $\bm{\theta}\sim\mathcal{P}$
$U(1)$-equivariant Computational basis measurement Givens Rotations [92] For all $\bm{\theta}$
$S_{n}$-equivariant Permutation invariant shadows [93] $\mathfrak{g}$-sim [94] For all $\bm{\theta}$
Matchgate circuit Expectation value of Pauli operators $\mathfrak{g}$-sim For all $\bm{\theta}$
Small angle initialization Pauli classical shadows Tensor Networks For $\bm{\theta}\sim\mathcal{P}$
Small Lie algebra $\mathfrak{g}$ Expectation value of algebra elements $\mathfrak{g}$-sim For all $\bm{\theta}$
Quantum generative modeling [^6] Here we refer to quantum circuit Born machines with the maximum mean discrepancy loss [30] and quantum generative adversarial networks with certain families of discriminators  [27] (both using shallow hardware efficient circuits). Not needed Tensor Networks For all $\bm{\theta}$
@@
Once we have identified $\mathcal{B}_{\lambda}$, we can proceed to study the adjoint action of $U(\bm{\theta})$ over the relevant $P_{\lambda}$ operator. The key insight here is to note that $U(\bm{\theta})P_{\lambda}U^{\dagger}(\bm{\theta})$ can only map to operators in $\mathcal{B}_{\lambda}$, and hence, can be somehow represented by its effective action on this subspace. The specific construction of this effective action is case dependent, but as detailed in the Appendix, we have been able to derive it for all considered barren plateau-free problems. In some cases, such as the aforementioned hardware efficient ansatz, one can trivially reduce $U(\bm{\theta})$ to a smaller dimensional unitary acting on $2L$ qubits by simply discarding all of the gates that are not in the light cone of the measurement (see Fig. 4(b)). In some other cases, one needs to employ classical simulation algorithms based on tensor networks, operator truncation, or Lie algebraic techniques [87, 96, 97, 94]. We refer the reader to the Appendix for additional details. It is worth noting that constructing the adjoint action of $U(\bm{\theta})$ on $\mathcal{B}_{\lambda}$ has a computational cost associated with it, which one needs to take into consideration to estimate the simulation cost. Importantly, while we have found algorithms whose cost is polynomial in the number of qubits, it can nevertheless have poor scaling.
The final task is to obtain the component of $\rho$ and $O$ in $\mathcal{B}_{\lambda}$. While it is possible to find the polynomially large subspaces and the effective adjoint of $U(\bm{\theta})$ fully classically, this will not be generically true for determining $O_{\lambda}$ and $\rho_{\lambda}$ (as we saw, for example, for the shallow HEA). While $O_{\lambda}$ is not much of a concern—since the classical description of $O$ could already contain this information—let us focus on obtaining $\rho_{\lambda}$. Recalling that, given an orthogonal basis $\{P_{j}\}_{j}$, one can write $\rho_{\lambda}=\sum_{P_{j}\in\mathcal{B}_{\lambda}}\left\langle\rho,P_{j}\right\rangle P_{j}/\sqrt{\left\langle P_{j},P_{j}\right\rangle}$, then we see that one needs to compute the expectation values $\left\langle P_{j},\rho\right\rangle$. While for simple states such as the all-zero state this task can be performed fully classically, for general input states $\rho$ this might not be the case.
@@
Problems in known polynomial subspaces are classically simulable (potentially requiring data from a quantum computer).
Consider a problem class such that $\mathcal{C}\subset\mathcal{C}_{{\rm polySub}}$. Then either
@@
$(ii.)\,\,\,\,\mathcal{C}\subset{\rm CSIM}_{\mathrm{QE}}\,\quad$ (14)
@@
In particular, this means one can simulate barren plateau-free loss functions $\mathcal{C}\subset\overline{{\rm BP}}$ living in a known polynomial subspace (as described in Claim 1) using a classical algorithm without the need to implement a parametrized quantum circuit on a quantum computer.
Again, for all barren plateau-free problems considered here, we have determined the relevant subspace, found efficient procedures to estimate the projections of the initial state and measurement operator, as well as for the action of the unitary. A summary of the resulting simulation algorithm and measurement protocols are reported in Table 2 and expanded upon in the Appendix.
@@
To wrap up, we highlight that combining Claim 1 and Claim 2 with our case-by-case analysis we can arrive at the conclusion: In none of the standard problem instances with provable absence of barren plateaus does the parametrized quantum circuit need to be implemented on a quantum computer in order to estimate the loss in polynomial time. Note that our results do not imply a dequantization of variational quantum computing as a whole, since a quantum device might still be needed for the initial data acquisition phase. Viewed this way, our results can be viewed positively as highlighting the potential of a different learning paradigm where the quantum computer is used non-adaptively to create a classical surrogate of the loss landscape. Still, in a way, Claims 1 and 2 do dequantize the variational part of the model and shed some serious doubts on the non-classicality of the information processing being done by a barren plateau-free loss function.

### v3

@@ III What leads to absence of barren plateaus?
If, however, the unitary $U(\bm{\theta})$ possesses additional structure, such as respecting some symmetry, then the loss can inherit this structure and potentially avoid the aforementioned issues. In particular, let us consider the adjoint action of $U(\bm{\theta})$ over the operator space $\mathcal{B}$, and let us analyze if it leads to either subspaces, or, effective subspaces. That is, given some $\bm{\theta}$ and operator $P$ in an appropriate orthogonal basis of $\mathcal{B}$ (such as a Pauli operator), we want to know: Where can $U(\bm{\theta})PU^{\dagger}(\bm{\theta})$ go? Note that one can, and should, also ask the question: Where can $U^{\dagger}(\bm{\theta})PU(\bm{\theta})$ go? But for simplicity of notation we will consider the first question.
Mathematically, this question can be answered by computing the inner products $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ for the rest of the basis elements $P_{j}$. For instance, if the inner product is non-zero for all $P_{j}$, then we know that the unitary can transform $P$ into an operator that can spread out and reach operators across all of $\mathcal{B}$. On the other hand, it could happen that the adjoint action of the unitary can only reach certain operators in $\mathcal{B}$. For instance, as shown in Fig. 3(a), the adjoint action of $U(\bm{\theta})$ can lead to well-defined subspaces in $\mathcal{B}$ such that $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ is only non-zero for operators in some $\mathcal{B}_{P}\subset\mathcal{B}$. This case can arise for instance in circuits with small Lie algebraic modules [22, 67, 103, 104, 52, 53, 54, 60] or in shallow-depth hardware efficient ansätze [24, 25, 26, 27].
A second case of interest arises when we have that $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ is non-zero for a wide range of operators, but is large only for operators in a given subspace $\mathcal{B}_{P}$. This case is depicted schematically in Fig. 3(b) and we will say that given $P$, the adjoint action of $U(\bm{\theta})$ leads to ‘effective’ subspaces, rather than proper ones. We note that effective subspaces can either arise for all $\bm{\theta}$ or with high probability for $\bm{\theta}\sim\mathcal{P}$. In the latter case, for most $\theta$ values for the $\left\langle U(\bm{\theta})PU^{\dagger}(\bm{\theta}),P_{j}\right\rangle$ are large only within a small subspace, but for some low probability $\bm{\theta}$ values, large overlaps outside this subspace could be observed. Effective subspaces appear, with high probability for $\bm{\theta}\sim\mathcal{P}$, for quantum convolutional neural networks [105, 15], in circuits with small angle initialization [39, 55, 61], or even due to the effects of noise [70].
@@
Shallow hardware efficient ansatz [24, 25, 26, 27, 29, 30, 31] $\mathcal{O}(\log(n))$ neighboring qubits (P) Local $O$, area law $\rho$
Generic shallow locally circuits [106, 107] [^1] See also the Supplemental Information for a detail proof of this claim. $\mathcal{O}(1)$-weight qubits (E) Local $O$, area law $\rho$
Quantum convolutional neural network [15] $\mathcal{O}(1)$-weight qubits (E) Local $O$, area law $\rho$
@@
$S_{n}$-equivariant [67] Permutation equivariant (P) $O\in\mathcal{B}_{\lambda}$, Tr[$\Tr[\rho_{\lambda}^{2}]\in\Omega(1/\operatorname{poly}(n))$
Matchgate circuit [54] Product of $\mathcal{O}(1)$ Majoranas (P) $O\in\mathcal{B}_{\lambda}$, Tr[$\Tr[\rho_{\lambda}^{2}]\in\Omega(1/\operatorname{poly}(n))$
Small angle initialization [39, 55, 61] $\mathcal{O}(\text{poly}(n))$ Pauli Operators (E) Local $O$, area law $\rho$
Small Lie algebra $\mathfrak{g}$ [52, 65, 58] Operators in $\mathfrak{g}$ (P) $O\in i\mathfrak{g}$, Tr[$\Tr[\rho_{\mathfrak{g}}^{2}]\in\Omega(1/\operatorname{poly}(n))$
Non-unital noisy-circuits [70] $\mathcal{O}(\log(n))$ qubits (E) Local $O$, any $\rho$
Dynamic circuits [73] $\mathcal{O}(1)$-weight qubits (E) Local $O$, any $\rho$
@@
Table 1: Problem classes in $\overline{{\rm BP}}$ and also in $\mathcal{C}_{{\rm polySub}}$. In this table, we present the problems considered in our analysis. In all cases, we also provide the references where absence of barren plateaus was proven, as well as the relevant polynomially large subspace $\mathcal{B}_{\lambda}$ (indicating whether it is proper (P) or effective (E)), and the states and measurement operators for which $\mathcal{C}$ is in $\overline{{\rm BP}}$. We refer the reader to Table 2 and the Supplemental Information for additional details on how the classical simulation is performed. Here ‘proj.’ is used as a short-hand for projector, while MPS stands for matrix product states [108].
To showcase the importance of having measurement operators and initial states which are “well-aligned” with the polynomial subspace, let us consider again the class of problems with shallow hardware efficient ansätze. We already saw that if $O$ is local (or contains relevant local terms), then $\mathcal{C}_{{\rm shallowHEA}}\subset\mathcal{C}_{{\rm polySub}}$. Thus, to know whether the loss will concentrate, we need to study the Hilbert-Schmidt norm of the projected $\rho$ into the subsets of $2L\in\mathcal{O}(\log(n))$ adjacent qubits in the backwards light cone of the local measurement. For instance, it is not hard to check that if $\rho$ is pure and follows a volume law of entanglement [24, 42], then $|\left\langle\rho_{\lambda},\rho_{\lambda}\right\rangle|$ will be exponentially close to zero, and the loss will be exponentially concentrated. On the other hand, if $\rho$ is pure and satisfies an area law of entanglement [109, 42], then the previous inner product will be, at most, polynomially vanishing. Putting these realizations together we see that
@@
[Standard provably barren plateau-free architectures live in classically identifiable polynomial subspaces.] For all standard problem classes $\mathcal{C}$ such that $\mathcal{C}\subset\overline{{\rm BP}}$, we have that
@@
By ‘standard problems’ we here refer to the conventional variational quantum architectures which have been proven to be in $\overline{{\rm BP}}$. In Table 1 we present a non-exhaustive list of such architectures. Therein, we report the relevant polynomial subspace, indicating whether it is proper or effective, as well as the conditions on the measurement operator and initial state necessary for absence of barren plateaus. We emphasize that in many cases, it was the proof of absence of barren plateau itself which allowed us to determine the reported information. The majority of strategies studied, including shallow hardware efficient ansatz and highly symmetrized models, lead to proper subspaces for all $\bm{\theta}$. The most important examples of models where one lives in effective subspaces (with high probability $\bm{\theta}\sim\mathcal{P}$) are quantum convolutional neural networks and small angle initialization strategies.
@@ IV Connection between absence of barren plateaus and simulability
$\mathcal{C}_{{\rm shallowHEA}}\subset{\rm CSIM}\,,\quad\text{or}\quad\mathcal{C}_{{\rm shallowHEA}}\subset{\rm QESIM}\,$, (12)
and it is clear that the loss can be estimated without ever needing to use a quantum computer to run the parametrized quantum circuit. Hence, no problem based on shallow hardware efficient ansätze with a local measurement can be outside of $\rm QESIM$}. We note that this result was originally reported in Ref. [30].It is worth additionally highlighting that, given the simplicity of the requisite measurement procedures here, one could lift the restriction that $\rho$ is preparable by a poly depth quantum circuit and also consider quantum states that result from quantum experiments running in polynomial time, e.g., from analog or non-unitary processes. This makes the quantum-enhanced classical simulation method more flexible than directly implementing the parameterized quantum circuit on hardware [110].
@@
Shallow hardware efficient ansatz Pauli classical shadows [95] Light-cone sim. reduced $U(\bm{\theta})$ $\forall\,\bm{\theta}$
Generic shallow locally circuits Pauli classical shadows Pauli Propagation [79] $\bm{\theta}\sim\mathcal{P}$
Quantum convolutional neural network [^3] In the Supplemental Information we also show that certain parameter restricted quantum convolutional neural networks can be simulated for all $\bm{\theta}$. For a general simulation of vanilla quantum convolutional neural networks see [78]. Pauli classical shadows Pauli Propagation $\bm{\theta}\sim\mathcal{P}$
$U(1)$-equivariant Computational basis measurement Givens Rotations [111] $\forall\,\bm{\theta}$
$S_{n}$-equivariant Permutation invariant shadows [112] $\mathfrak{g}$-sim [113] $\forall\,\bm{\theta}$
Matchgate circuit Expectation value of Pauli operators $\mathfrak{g}$-sim $\forall\,\bm{\theta}$
Small angle initialization Pauli measurements Tensor Networks [82], Pauli Prop. [80] $\bm{\theta}\sim\mathcal{P}$
Small Lie algebra $\mathfrak{g}$ Expectation value of algebra elements $\mathfrak{g}$-sim $\forall\,\bm{\theta}$
Non-unital noisy-circuits Not needed Light-cone sim. [70], Pauli Prop. [114] $\bm{\theta}\sim\mathcal{P}$
Dynamic circuits Pauli classical shadows Pauli Propagation [79] $\bm{\theta}\sim\mathcal{P}$
Quantum generative modeling [^4] Here we refer to quantum circuit Born machines with the maximum mean discrepancy loss [32] and quantum generative adversarial networks with certain families of discriminators  [29] (both using shallow hardware efficient circuits). Not needed Tensor Networks $\forall\,\bm{\theta}$
@@
Once we have identified $\mathcal{B}_{\lambda}$, we can proceed to study the adjoint action of $U(\bm{\theta})$ over the relevant $P_{\lambda}$ operator. The key insight here is to note that $U(\bm{\theta})P_{\lambda}U^{\dagger}(\bm{\theta})$ can only map to operators in $\mathcal{B}_{\lambda}$, and hence, can be somehow represented by its effective action on this subspace. The specific construction of this effective action is case dependent, but as detailed in the Supplemental Information (which contains additional details and proofs, which contains Refs. [120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159].), we have been able to derive it for all considered barren plateau-free problems. In some cases, such as the aforementioned hardware efficient ansatz, one can trivially reduce $U(\bm{\theta})$ to a smaller dimensional unitary acting on $2L$ qubits by simply discarding all of the gates that are not in the light cone of the measurement (see Fig. 4(b)). In some other cases, one needs to employ classical simulation algorithms based on tensor networks, operator truncation, or Lie algebraic techniques [104, 116, 117, 113]. We refer the reader to the Supplemental Information for additional details. It is worth noting that constructing the adjoint action of $U(\bm{\theta})$ on $\mathcal{B}_{\lambda}$ has a computational cost associated with it, which one needs to take into consideration to estimate the simulation cost. Importantly, while we have found algorithms whose cost is polynomial in the number of qubits, it can nevertheless have poor scaling.
The final task is to obtain the component of $\rho$ and $O$ in $\mathcal{B}_{\lambda}$, which we respectively denote as $\rho_{\lambda}$ and $O_{\lambda}$. While it is possible to find the polynomially large subspaces and the effective adjoint of $U(\bm{\theta})$ fully classically, this will not be generically true for determining $O_{\lambda}$ and $\rho_{\lambda}$ (as we saw, for example, for the shallow HEA). While $O_{\lambda}$ is not much of a concern—since the classical description of $O$ could already contain this information—let us focus on obtaining $\rho_{\lambda}$. Recalling that, given an orthogonal basis $\{P_{j}\}_{j}$, one can write $\rho_{\lambda}=\sum_{P_{j}\in\mathcal{B}_{\lambda}}\left\langle\rho,P_{j}\right\rangle P_{j}/\sqrt{\left\langle P_{j},P_{j}\right\rangle}$, then we see that one needs to compute the expectation values $\left\langle P_{j},\rho\right\rangle$. While for simple states such as the all-zero state this task can be performed fully classically, for general input states $\rho$ this might not be the case.
@@
[Problems in known polynomial subspaces are classically simulable (potentially requiring data from a quantum computer).] Consider a problem class such that $\mathcal{C}\subset\mathcal{C}_{{\rm polySub}}$. Then either
@@
$(ii.)\,\,\,\,\mathcal{C}\subset{\rm QESIM}\,\quad$ (14)
@@
In particular, this opens the door towards the simulation of barren plateau-free loss functions $\mathcal{C}\subset\overline{{\rm BP}}$ living in a known polynomial subspace (as described in Claim 1) using a classical algorithm without the need to implement a parametrized quantum circuit on a quantum computer.
Again, for all barren plateau-free problems considered here, we have determined the relevant subspace, found efficient procedures to estimate the projections of the initial state and measurement operator, as well as for the action of the unitary. A summary of the resulting simulation algorithm and measurement protocols are reported in Table 2 and expanded upon in the Supplemental Information.
@@
To wrap up, we highlight that combining Claim 1 and Claim 2 with our case-by-case analysis we can arrive at the conclusion: For the families of problems and architectures with provable absence of barren plateaus we analyzed, it does not seem that the parametrized quantum circuit need to be implemented on a quantum computer in order to estimate the loss in polynomial time. Note that our results do not imply a dequantization of variational quantum computing as a whole, since a quantum device might still be needed for the initial data acquisition phase. Viewed this way, our results can be viewed positively as highlighting the potential of a different learning paradigm where the quantum computer is used non-adaptively to create a classical surrogate of the loss landscape. Still, in a way, Claims 1 and 2 do, in a sense, dequantize the variational part of the model and shed some serious doubts on the non-classicality of the information processing being done by a barren plateau-free loss function.